### v3

# Diagnostic-Guided Dynamic Profile Optimization for LLM-based User Simulators in Sequential Recommendation

Abstract
Recent advances in large language models (LLMs) have enabled realistic user simulators for developing and evaluating recommender systems (RSs). However, existing LLM-based simulators for RSs face two major limitations: (1) static and single-step prompt-based inference that leads to inaccurate and incomplete user profile construction; (2) unrealistic and single-round recommendation-feedback interaction pattern that fails to capture real-world scenarios. To address these limitations, we propose DGDPO (Diagnostic-Guided Dynamic Profile Optimization), a novel framework that constructs user profile through a dynamic and iterative optimization process to enhance the simulation fidelity. Specifically, DGDPO incorporates two core modules within each optimization loop: firstly, a specialized LLM-based diagnostic module, calibrated through our novel training strategy, accurately identifies specific defects in the user profile. Subsequently, a generalized LLM-based treatment module analyzes the diagnosed defect and generates targeted suggestions to refine the profile. Furthermore, unlike existing LLM-based user simulators that are limited to single-round interactions, we are the first to integrate DGDPO with sequential recommenders, enabling a bidirectional evolution where user profiles and recommendation strategies adapt to each other over multi-round interactions. Extensive experiments conducted on three real-world datasets demonstrate the effectiveness of our proposed framework. 00footnotetext: denotes the corresponding authors.
Introduction
User simulators (Balog and Zhai 2023), which aim to mimic user behaviors and decision-making patterns, are crucial for advancing recommender systems (RSs) (Luo et al. 2022). They offer a controllable and cost-effective paradigm for developing and evaluating RSs, reducing the reliance on real-user experiments that are often expensive and ethically sensitive. While early user simulators ranged from rule-based (Rohde et al. 2018) to deep learning-based approaches (Shi et al. 2019b; Zhao et al. 2021), these traditional methods fail to explicitly represent users’ complex semantic preferences and generate diverse user interaction behaviors. To address these issues, Large Language Models (LLMs) have recently emerged as a new paradigm (Sun et al. 2024b). With their advanced cognitive and generative capabilities, LLMs are typically integrated into an agent framework (Park et al. 2023) that incorporates components such as Profile, Memory, and Action to emulate user interactions with RSs. For example, Agent4Rec (Zhang et al. 2024a) and RecAgent (Wang et al. 2025) prompt LLMs to generate user profile, enabling them to simulate diverse behaviors that mirror the complexity of real users.
However, existing LLM-based user simulators face two major limitations: (1) the inaccuracy and incompleteness of profile construction due to the static and single-step prompt-based inference. By generating the user profile only at initialization, user simulators produce a static representation that cannot adapt to the evolving user interests across multi-round interactions. Thus, any inaccuracy or incompleteness in this initial profile persists uncorrected, causing the simulated behavior to progressively diverge from real user actions throughout the simulation; (2) the unrealistic and single-round recommendation-feedback interaction pattern. Most existing simulators are typically integrated with static recommenders, e.g., MF (Koren, Bell, and Volinsky 2009), failing to simulate authentic, multi-round interactions where user profile and RS strategies should adapt and evolve based on the latest feedback.
ML-1MBooksMovies & TV02020404060608080100100Accuracy (%) Prompt-based general-purpose LLM (GPT-4o-mini) Specialized LLM-based diagnostic module (Qwen3-0.6B)
Figure 1: Accuracy comparison for profile defect identification. Our specialized LLM-based diagnostic module significantly outperforms the prompt-based general-purpose LLM. Details are included in Appendix D.
Therefore, we propose a novel Diagnostic-Guided Dynamic Profile Optimization (DGDPO) framework for a more realistic simulation. However, developing this framework faces two primary challenges. (CH1): How to build a reliable mechanism for dynamic profile construction? While existing methods (Shu et al. 2024) like self-reflection (Sun et al. 2024a) attempt iterative refinement, they often rely on general-purpose LLMs, thus may not be reliable due to potential hallucinations. As shown in Fig. 1, directly prompting general-purpose LLMs only achieves an average accuracy of 62.78% on profile defect identification. This highlights the necessity of a more targeted strategy that can accurately diagnose the specific type of defect, such as inaccurate and/or incomplete, and then apply a tailored refinement. To achieve this, we first design a specialized LLM-based diagnostic module through domain-adaptive pre-training and defect-specific fine-tuning, enabling reliable identification of profile defects. As shown in Fig. 1, the module achieves an average accuracy of 92.20%, significantly outperforming the general-purpose LLM. We then design a generalized LLM-based treatment module that uses this precise diagnosis to analyze the defect and generate targeted refinements for the profile. Instead of single-step prompting, our framework iteratively optimizes the profile by traversing the user’s interaction history step-by-step, to dynamically enhance its accuracy and comprehensiveness. (CH2): How to achieve realistic multi-round interactions where user profiles and recommender strategies mutually influence and evolve over time? Inspired by the multi-round interaction nature of real-world recommendation scenarios, we propose to incorporate the user simulator with sequential recommenders (SRs) (Sun et al. 2025). This establishes a genuine bidirectional evolution: the user profile is updated based on interactions, while the SRs adaptively adjust their recommendation strategies in response to the evolving user behavior.
Our main contributions are three-fold: (1) To address CH1, we propose a novel framework DGDPO, which integrates a specialized LLM-based diagnostic module and a generalized LLM-based treatment module. By iteratively optimizing the user profile through step-by-step traversal of user interactions, the framework dynamically refines the user profile, resulting in a more accurate and comprehensive representation. (2) To address CH2, we establish a realistic and multi-round interaction between the user simulator and sequential recommenders. This enables a bidirectional evolution of user profiles and recommendation strategies, offering a more credible evaluation environment. (3) We conduct extensive experiments on three real-world datasets, empirically validating the effectiveness of our DGDPO framework.
Related Works
Traditional User Simulators for RSs. Early user simulators for RSs progressed from rule-based approaches (Rohde et al. 2018; Ie et al. 2019; Shi et al. 2019a) to advanced data-driven techniques (Huang et al. 2020), including methods based on reinforcement learning (RL) (Shi et al. 2019b; Zou et al. 2020), generative adversarial networks (GANs) (Chen et al. 2019; Shi et al. 2019b; Zhao et al. 2021), and transformers (Zhao et al. 2023; Afzali et al. 2023). However, they lack external context and knowledge, as well as the sophisticated cognitive reasoning mechanism that real users leverage during decision-making processes.
LLM-based User Simulators for RSs. The advent of LLMs has introduced novel paradigms for user simulation in RSs (Cai et al. 2025; Ma et al. 2025). LLM-based simulators have been widely adopted across various applications, including evaluating multi-round conversational systems (Friedman et al. 2023; Zhu, Huang, and Sang 2024) and generating high-quality interaction data for RL-based algorithms (Corecco et al. 2024; Ebrat, Paradalis, and Rueda 2024; Zhang et al. 2025). Current works on user profile construction can be categorized into two types. Some works (Zhang et al. 2024a; Wang et al. 2025) create user profile once at initialization, resulting in static representations that cannot capture evolving user preferences. Others employ self-reflection (Zhang et al. 2024b; Shu et al. 2024) for iterative refinement. However, they suffer from two limitations: (1) they adopt general-purpose LLMs which may generate unreliable profile due to potential hallucinations; (2) they rely on unrealistic single-round interactions, lacking dynamic adaptation to user feedback and failing to simulate the bidirectional evolution between user profile and recommendation strategies in realistic interactions.
Sequential Recommenders. SRs (Du et al. 2024) have evolved from early approaches combining Markov Chains with matrix factorization (Rendle, Freudenthaler, and Schmidt-Thieme 2010), to deep learning models. These include CNNs for local pattern extraction (Tang and Wang 2018), RNNs to capture temporal dependencies (Hidasi 2015; Huang et al. 2025), attention mechanisms to focus on relevant historical items (Li et al. 2017; Kang and McAuley 2018; Liu et al. 2018; Sun et al. 2019), and GNNs to model high-order relationships within sequential data (Wu et al. 2022; Yu et al. 2024; Li et al. 2024; Liu, Xia, and Huang 2024). Despite their success, they have primarily been evaluated in a static mode, which may fail to fully reflect the dynamic complexity of real-world user interactions.
The Proposed Method
Framework Overview. DGDPO consists of two core modules that iteratively optimize the user profile in a dynamic manner, as shown in Fig. 2. First, a specialized LLM-based diagnostic module identifies profile defects by detecting discrepancies between simulated and real user behavior. Then, a generalized LLM-based treatment module generates targeted refinements to optimize the profile based on the diagnosis. This optimization process is executed iteratively across a user’s historical interaction sequence (i.e., step-by-step traversal of user interactions). For a more realistic simulation, DGDPO is further incorporated with SRs to establish multi-round interactions, enabling bidirectional evolution of user profiles and recommendation strategies.
Figure 2: The overall framework of our proposed DGDPO.
The DGDPO Framework (CH1)
Given a user’s historical interaction sequence $\mathcal{D}_{u}=[i_{1},i_{2},...,i_{T}]$ ordered by timestamp, we split it into two portions: $\mathcal{D}_{u}^{(ini)}=[i_{1},i_{2},...,i_{n}]$, and $\mathcal{D}_{u}^{(opt)}=[i_{n+1},...,i_{T}]$. For effective and efficient user profile construction, we first create an initial user profile $\mathcal{P}_{u}^{(ini)}$ for each simulated user $u$. It is derived from $\mathcal{D}_{u}^{(ini)}$ via a single-step prompt-based inference method (Zhang et al. 2024a), serving as the foundation user profile. Then, we leverage $\mathcal{D}_{u}^{(opt)}$ to dynamically optimize $\mathcal{P}_{u}^{(ini)}$ via our proposed DGDPO, ultimately obtaining an optimized user profile $\mathcal{P}_{u}^{(opt)}$.
This optimization process is triggered by collecting discrepancy cases where the simulated behavior $\mathcal{B}_{u}^{(\mathrm{sim})}$ of user $u$ mismatches his real behavior $\mathcal{B}_{u}^{(\mathrm{real})}$ observed in the datasets (i.e., the simulator declines to interact with an item that the real user has interacted with). Specifically, we present the first item $i_{n+1}\in\mathcal{D}_{u}^{(opt)}$ to the user simulator built upon $\mathcal{P}_{u}^{(ini)}$. Then, the user simulator determines whether to interact (i.e., click) with $i_{n+1}$ and produces an explanatory reason for the decision. We traverse items in $\mathcal{D}_{u}^{(opt)}$ to collect a batch ($N_{1}$) of such discrepancy cases, which then serve as the input for our optimization workflow.
The Specialized LLM-based Diagnostic Module.
Given a discrepancy case, this module aims to accurately identify the defect of the user profile via a specialized LLM. As such, it overcomes the inherent unreliability of general-purpose LLMs for such specific diagnostic tasks, providing more accurate signals to guide the optimization process.
First, we categorize each profile defect into one of three types based on its root cause: (1) Inaccurate: the profile contains preference descriptions that contradict the real user’s observed behavior. (2) Incomplete: the profile lacks the specific preference descriptions required to explain the user’s interaction with a given item. (3) Inaccurate & Incomplete: the profile exhibits both of the aforementioned issues. Then, we formally define this diagnostic task as a function $f_{\text{diag}}$. Given a user profile $\mathcal{P}_{u}$ and the corresponding discrepancy case $\mathcal{C}_{j}$, it outputs the corresponding defect type $l=f_{\text{diag}}(\mathcal{P}_{u},\mathcal{C}_{j})$ from the three defined types.
The key challenge now is how to accurately predict $l$. A straightforward way is to directly prompt general-purpose LLMs. However, their performance may be suboptimal due to limited domain knowledge and inherent hallucination (as illustrated in Fig. 1). Therefore, we employ a delicately designed training strategy (i.e., domain-adaptive pre-training and defect-specific fine-tuning) to build a specialized LLM, transforming it into a specialized expert in profile defect identification by injecting extensive domain knowledge.
Domain-Adaptive Pre-training. This stage is to inject the model with the underlying semantics and reasoning patterns in recommendation. By pre-training on a large corpus of simulated interactions, the model learns the complex relationships between user profiles, item attributes, and interaction decisions, providing a robust foundation for the downstream defect identification task.
To construct the pre-training corpus, for each user $u$, we leverage his initial sequence data $\mathcal{D}_{u}^{(ini)}$. For each item $i_{t}\in\mathcal{D}_{u}^{(ini)}$, we instruct the user simulator with profile $\mathcal{P}_{u}^{(ini)}$ to make a binary interaction decision for $i_{t}$ and generate a textual rationale. To ensure the data quality, we only consider the case where the user simulator’s decision matches with the real user $u$’s behavior (i.e., choosing to interact with $i_{t}$) and user $u$ gave item $i_{t}$ a high rating ($\geq$ 3 out of 5). Then, the user simulator’s output, together with $\mathcal{P}_{u}^{(ini)}$ and relevant information (e.g., title, genres, and rating) of $i_{t}$, is used to construct a comprehensive corpus of structured texts, where the details are shown in Appendix A.
Given the constructed corpus, we perform domain-adaptive pre-training on a small-scale general-purpose LLM using a language modeling objective, where the goal is to predict the next token in a sequence. This encourages the LLM to internalize the domain-specific logical flow of profile-based user simulation, defined as:
$\mathcal{L}_{\text{pretrain}}=-\textstyle\sum_{{\scriptstyle t=1}}^{{\scriptstyle T_{Pre}}}\log P(x_{t}|x_{<t},\theta),$ (1)
where $x=(x_{1},\ldots,x_{T_{Pre}})$ represents a text sequence from the corpus with length $T_{Pre}$, and $\theta$ are the model parameters.
Defect-Specific Fine-tuning. Next, we further fine-tune the LLM for the specific profile defect diagnosis task. A primary challenge lies in the absence of available, human-annotated data for defect identification. To overcome this, we develop a synthetic data generation pipeline that creates a high-quality, labeled corpus for fine-tuning purposes.
The pipeline starts with the high-quality, non-defective profile-item pairs that constitute our pre-training corpus. Building upon these non-defective pairs, we subsequently perform a series of targeted modifications to generate synthetic user profiles with different defect types: (1) Inaccurate: to create a contradiction, we first identify preference descriptions in the profile relevant to the target item (e.g., its genre). We then generate a new description with the opposite sentiment (e.g., transforming “enjoys comedy films” to “dislikes comedy films”) and use it to replace the original statement. (2) Incomplete: to simulate a lack of information, we randomly remove a proportion of descriptions relevant to the target item from the profile, making the profile insufficient to explain the user’s interaction. (3) Inaccurate & Incomplete: we first apply the “incomplete” modification, followed by the “inaccurate” modification. Using these synthetic profiles, we then format the data for instruction fine-tuning, where each sample has four parts: System, Instruction, Input, and Output, as shown in Appendix A.
To fine-tune the model, we employ a next token prediction loss objective with a modification: the loss is only calculated on the tokens in the Output field. The tokens from the other fields are masked from the loss calculation, forcing the model to focus exclusively on more accurate profile defect diagnosis. Thus, the fine-tuning objective is to minimize the loss only over the target response tokens:
$\mathcal{L}_{\text{finetune}}=-\sum\nolimits_{t\in T_{\text{Output}}}\log P(x_{t}|x_{<t},\theta),$ (2)
where $T_{\text{Output}}$ is the set of token indices corresponding to the Output field, and $\theta$ are the model parameters.
In summary, our delicately designed training strategy transforms the general-purpose LLM into a highly accurate specialized diagnostic module. It provides a trustworthy and structured diagnostic signal that is essential for guiding the subsequent treatment module.
Input: $\mathcal{D}_{u}^{(ini)}$, $\mathcal{D}_{u}^{(opt)}$, $N_{1}$
Output: $\mathcal{P}_{u}^{(opt)}$
$\mathcal{C}$ $\leftarrow$ [ ];
  // Initialize discrepancy cases list
[…]
  // Single-step prompt-based inference
1 for item $i$ in $\mathcal{D}_{u}^{(opt)}$ do
[…]
    // Generate simulated behavior
[…]
    // Get real behavior
2 if $\mathcal{B}_{u}^{(\mathrm{sim})}$ $\neq$ $\mathcal{B}_{u}^{(\mathrm{real})}$ then
3    $\mathcal{C}.append(i,\mathcal{B}_{u}^{(\mathrm{sim})}$, $\mathcal{B}_{u}^{(\mathrm{real})}$);
4    if $|\mathcal{C}|=N_{1}$ then
         // The Specialized LLM-based Diagnostic Module
5       $\{l\}\leftarrow f_{\text{diag}}(\mathcal{P}_{u}^{(ini)},\mathcal{C})$;
         // The Generalized LLM-based Treatment Module
       $\{r\}\leftarrow\text{Infer}(\mathcal{P}_{u}^{(ini)},\mathcal{C},\{l\})$;
          // Infer defect reasons
         // Generate optimization suggestions
6       $\mathcal{O}\leftarrow\text{GenOpt}(\mathcal{P}_{u}^{(ini)},\mathcal{C},\{l\},\{r\})$;
       $\tilde{\mathcal{P}}_{u}^{(ini)}\leftarrow\text{Refine}(\mathcal{P}_{u}^{(ini)},\mathcal{O})$;
          // Refine the profile
7       $\mathcal{C}$ $\leftarrow$ [ ];
8       $\mathcal{P}_{u}^{(ini)}$ $\leftarrow$ $\tilde{\mathcal{P}}_{u}^{(ini)}$;
9
10
[…]
12 return ${\mathcal{P}}_{u}^{(opt)}$;
Algorithm 1 DGDPO
The Generalized LLM-based Treatment Module.
This module is responsible for the complex reasoning and modification tasks. While the specialized diagnostic module excels at identifying what defects exist within a profile, understanding why these defects occur and determining how to address them effectively requires sophisticated common-sense reasoning and high-fidelity generative capabilities. Consequently, we employ a powerful general-purpose LLM, leveraging its superior analytical reasoning and text generation capabilities to carry out the following treatment process [^1] Detailed prompts are attached in Appendix B..
Specifically, we first prompt the LLM by providing it with the current user profile $\mathcal{P}_{u}^{(ini)}$, the target item $i$, and defect label $l$, to generate a comprehensive textual explanation (i.e., inferred reason $r$) towards the diagnosis made by the diagnosis module. Given the inferred reason, we further instruct the LLM to generate a set of concrete optimization suggestions (i.e., targeted modifications) denoted as $\mathcal{O}$. For instance, an inaccurate defect may result in a suggestion to correct contradictory preferences, while an incomplete defect would lead to a suggestion to add missing details. Finally, we prompt the LLM to apply the suggestions $\mathcal{O}$ to produce an updated user profile $\tilde{\mathcal{P}}_{u}^{(ini)}$. In this step, specific constraints are enforced to (1) maintain the profile’s overall coherence and descriptive style, and (2) avoid mentioning specific item details, ensuring the resulting profile captures general user preferences.
Iterative Optimization.
Given the updated profile $\tilde{\mathcal{P}}_{u}^{(ini)}$, our framework continues to traverse the remaining interactions in $\mathcal{D}_{u}^{(opt)}$ to collect the next batch of discrepancy cases. Once a batch is accumulated, the complete workflow is executed: the diagnostic module provides a diagnosis for each case, which subsequently guides the treatment module to produce the next updated profile. This process repeats until all interactions in $\mathcal{D}_{u}^{(opt)}$ are processed, ultimately yielding the final optimized user profile $\mathcal{P}_{u}^{(opt)}$. The complete process is detailed in Algorithm 1.
Bidirectional Evolution between User Profiles and Recommendation Strategies (CH2)
Existing LLM-based user simulators are primarily integrated with static recommenders in page-by-page scenarios (Zhang et al. 2024a), which essentially represent a one-time recommendation result split into multiple pages. This fails to capture realistic scenarios where users continuously interact with the RSs, leading to the dynamic evolution of both user profiles and recommendation strategies through user feedback and interactions. To enable realistic multi-round interactions, we propose to incorporate our optimized user simulator DGDPO with sequential recommenders (SRs). Unlike traditional static models, SRs excel at capturing temporal dynamics, which makes them ideal for modeling the bidirectional evolution of user preferences and recommendation strategies over time.
Specifically, in the multi-round interaction scenario, our DGDPO and SRs engage in a continuous interaction process that consists of four steps: (1) Based on the user’s interaction history, the SRs generate a ranked list of candidate items $\mathcal{X}$ for the optimized simulator with profile $\mathcal{P}_{u}^{(opt)}$, where one item is the ground truth, and the others are randomly sampled negative items; (2) The user simulator decides whether to interact with the recommended items, where at most one item can be selected for interaction in each round; (3) When an item is selected, the simulator performs an update to $\mathcal{P}_{u}^{(opt)}$ to reflect the recent interaction pattern. In particular, we design two strategies for profile update: (i) Update without Ground Truth (w/o GT): the profile is updated using the item that the simulator chooses to interact with. (ii) Update with Ground Truth (w/ GT): the profile is updated using the ground truth positive item to maintain alignment with real user behavior. In both strategies, the profile update is performed via a carefully controlled prompt that instructs the LLM to make minimal, incremental adjustments while preserving the profile’s overall coherence; and (4) Based on this newly observed interaction, the SRs update their internal hidden state for next recommendation. This loop continues until the simulator either chooses to exit or reaches the maximum rounds of interactions.
Experiments and Results
We conduct extensive experiments to answer five research questions. (RQ1) Can DGDPO improve the fidelity of user simulators? (RQ2) How do different profile update strategies and SRs impact the simulation fidelity in multi-round interaction scenarios? (RQ3) How do various components of DGDPO affect its performance? (RQ4) How do essential hyper-parameters affect DGDPO? (RQ5) How does DGDPO construct user profile with its optimization process?
ML-1M Books Movies & TV
#Users 6,040 208,864 23,969
#Items 3,416 241,725 25,830
#Interactions 999,611 10,865,527 1,053,194
Avg. Sequence Length 198.41 49.16 48.59
Sparsity 95.16% 99.98% 99.83%
Table 1: Dataset statistics after preprocessing.
Datasets ML-1M Books Movies & TV
Methods Prec Recall Acc F1 Prec Recall Accuracy F1 Prec Recall Acc F1
Traditional KuaiSim 0.1067 0.6304 0.6651 0.1825 0.1423 0.6107 0.6728 0.2308 0.0832 0.5917 0.6228 0.1459
LLM-based Agent4Rec 0.1031 0.6250 0.6600 0.1770 0.1521 0.6431 0.6920 0.2460 0.0823 0.6013 0.6322 0.1448
LLM-US 0.0978 0.6208 0.6702* 0.1690 0.1457 0.6243 0.6832 0.2363 0.0765 0.5984 0.6281 0.1357
RAH 0.1104 0.6413 0.6557 0.1884 0.1565 0.6348 0.6874 0.2511 0.0854 0.6078 0.6359 0.1498
AgentCF 0.1157* 0.6424* 0.6652 0.1961* 0.1598* 0.6581* 0.6953* 0.2572* 0.0871* 0.6105* 0.6394* 0.1524*
DGDPO 0.1430 0.7365 0.7174 0.2395 0.2058 0.7556 0.7570 0.3235 0.1038 0.6913 0.6884 0.1805
Improve 23.60% 14.65% 7.04% 22.13% 28.78% 14.81% 8.87% 25.82% 19.17% 13.24% 7.66% 18.44%
Table 2: Performance comparison on all datasets, where the best and runner-up results are highlighted in bold and marked by ‘*’; and ‘Improve’ indicates the relative improvements comparing the best and runner-up results. Statistical significance of the improvement is determined by a paired t-test with $p$-value << 0.01.
11223344556677889910100.050.050.0550.0550.060.06PrecisionFPMC
11223344556677889910100.080.080.090.090.10.1PrecisionNARM
11223344556677889910100.10.10.1050.1050.110.11PrecisionSASRec
11223344556677889910100.1050.1050.1150.1150.1250.125PrecisionLRURecw/o UpdateUpdate w/o GTUpdate w/ GT
Figure 3: The impact of profile update strategies on DGDPO in multi-round interactions on Books dataset.
Experimental Setup
Datasets.
We use three real-world datasets from various domains. Specifically, ML-1M (Harper and Konstan 2015) is widely used in movie recommendations. Amazon Books (Books) and Amazon Movies & TV (Movies & TV) are product review datasets collected from Amazon for Books and Movies & TV categories (Ni, Li, and McAuley 2019), respectively. We filter out users and items with less than 5 interactions in ML-1M, and those with less than 20 interactions in Amazon datasets. To support multi-round interactions, we first sort each user’s interaction chronologically. We set the maximum sequence length to 200 for ML-1M and 50 for Amazon datasets. For each user $u$, we select the most recent 10 items as the test set, the 11th most recent item as the validation set, and the remaining as the training set (i.e., $\mathcal{D}_{u}$). Particularly, we further split $\mathcal{D}_{u}$ into $\mathcal{D}_{u}^{(ini)}$ and $\mathcal{D}_{u}^{(opt)}$ with the ratio of $\alpha:(1-\alpha)$. The statistics of these datasets after preprocessing are summarized in Table 1.
Baselines.
We compare DGDPO with five baselines for user profile construction in two categories. In traditional methods, KuaiSim (Zhao et al. 2023) is a transformer-based model for user profile construction. In LLM-based methods, Agent4Rec (Zhang et al. 2024a) utilizes a single-step prompt-based inference. LLM-US (Zhang et al. 2025) generates explanations of user ratings as profile. RAH (Shu et al. 2024) and AgentCF (Zhang et al. 2024b) employ self-reflection and collaborative reflection, respectively. To enable multi-round interactions, we integrate the user simulator with four representative SRs. The conventional method FPMC (Rendle, Freudenthaler, and Schmidt-Thieme 2010) combines matrix factorization with first-order Markov chains. The deep learning-based methods include NARM (Li et al. 2017), an RNN-based model with an attention mechanism; SASRec (Kang and McAuley 2018), using unidirectional self-attention; and LRURec (Yue et al. 2024), leveraging a linear recurrent unit-based model.
Implementation Details.
Following state-of-the-arts (Shu et al. 2024; Zhang et al. 2024a), we randomly sample 1000 users from each dataset to serve as simulated users. Our diagnostic module is based on lightweight Qwen3-0.6B (Yang et al. 2025), while the treatment module employs gpt-4o-mini via its API with a temperature of 0. All LLM-based baselines also utilize GPT-4o-mini as their backbone. For DGDPO, we set $N_{1}=4$, $\alpha=0.6$, and $\mathcal{X}=20$ for all datasets. We build upon Agent4Rec (Zhang et al. 2024a) as the foundation user simulator and conduct ten-round interactions for each user. For KuaiSim and SRs methods, we tune them by Optuna (optuna.org) with 50 trials and train for up to 100 epochs with an early stopping mechanism (Sun et al. 2022). All experiments are conducted on an NVIDIA V100 GPU. More detailed implementations are in Appendix C.
Evaluation Metrics.
Following (Luo et al. 2022; Zhang et al. 2024a), we adopt Precision (Prec), Recall, Accuracy (Acc), and F1 Score (F1) as the evaluation metrics to measure the consistency between the simulated and real users.
Results and Analysis
Overall Comparison (RQ1).
Table 2 presents the performance comparison of all methods. The results show that DGDPO achieves the best performance across all datasets and metrics, with average relative improvements of 16.87%, 18.82%, and 14.63% over the runner-up methods across the three datasets, respectively. These substantial improvements demonstrate the effectiveness of our method. In addition, we observe that DGDPO provides a particularly substantial boost in Precision, with relative improvements of 23.60%, 28.78%, and 19.17% across the three datasets. This indicates that DGDPO can accurately identify and guide the optimization of defective user profiles, generating more precise and faithful user representations that prevent incorrect item selections and significantly enhance simulation precision.
11223344556677889910100.0490.0490.0510.0510.0530.0530.0550.055PrecisionFPMC
11223344556677889910100.070.070.0750.0750.080.08PrecisionNARM
11223344556677889910100.090.090.0950.0950.10.1PrecisionSASRec
11223344556677889910100.0980.0980.1020.1020.1060.106PrecisionLRURecAgentCFDGDPOKuaiSim
Figure 4: The performance of different user simulators in multi-round interactions on Books dataset.
Analysis of Multi-round Interactions (RQ2).
Fig. 3 depicts the performance of DGDPO across ten interaction rounds on Books dataset under different profile update strategies [^2] Hereafter, we only show the results on Books due to space limitation, and similar trends can be noted on the other two datasets.. The results unveils that dynamic profile update in multi-round interactions can further improve the simualtion accuracy. (1) For ‘w/o Update’, the performance generally exhibits minimal fluctuations across multi-round interactions and is the worst among all strategies. With profile unchanged, ‘w/o Update’ fails to capture the evolution of user preference during multi-round interactions, thereby limiting the simulator’s long-term fidelity. (2)Update w/o GT’ exhibits performance divergence when interacting with different SRs. Specifically, with FPMC, the performance shows a declining trend. In contrast, with advanced deep learning-based SRs, we observe steady performance improvements. This occurs because advanced SRs rank ground truth items higher, and LLM-based user simulators, sensitive to item positions (Hou et al. 2024), are more likely to interact with them when they are prominently placed, resulting in more accurate simulation. (3)Update w/ GT’ performs the best among all methods. This is attributed to an idealized update strategy that always uses the ground truth item to update the user profile, regardless of the simulator’s actions. However, in practical applications, ground truth items are often unavailable, which limits the applicability of this strategy. Therefore, it underscores the importance of developing more advanced user simulators that closely replicate real user behaviors, enabling more effective and reliable updates to the user profile during long-term interactions.
To further validate the effectiveness of our DGDPO in the multi-round interactions, we compare it against representative baselines. All simulators adopt the ‘w/o Update’ strategy and interact with four different SRs. Fig. 4 presents the results. First, DGDPO consistently achieves the highest Precision across every interaction round, indicating that it better captures user preferences and yields more realistic simulation. Second, although AgentCF underperforms DGDPO, it significantly outperforms traditional KuaiSim, reinforcing the effectiveness of LLM-based user simulators.
Ablation Study (RQ3).
Variant Prec Recall Acc F1
w/o iteration 0.1702 0.7204 0.7262 0.2753
w/o specialized LLM 0.1711 0.7219 0.7348 0.2760
w/o fine-tuning 0.1803 0.7220 0.7404 0.2885
w/o pre-training 0.2027 0.7411 0.7419 0.3183
DGDPO 0.2058 0.7556 0.7570 0.3235
Table 3: Ablation study of DGDPO on Books dataset.
Table 3 presents the results of ablation study. First, DGDPO consistently outperforms its ‘w/o iteration’ variant, which conducts optimization only once without iteratively traversing interactions step-by-step for dynamic refinement. This showcases the importance of iterative optimization for progressively enhancing user profile quality and achieving superior fidelity. Second, we validate the effectiveness of our diagnostic module by comparing DGDPO with the ‘w/o specialized LLM’ variant, where the specialized LLM is replaced with a general-purpose LLM. The results confirm that domain-specific specialized LLM enables more reliable and precise diagnostic guidance. Third, by removing either domain-adaptive pre-training from the specialized LLM (‘w/o pre-training’) or defect-specific fine-tuning (‘w/o fine-tuning’), the performance significantly drops compared to DGDPO, validating the usefulness of our designed training strategy. Notably, removing fine-tuning results in more substantial performance degradation, indicating a more critical role of defect-specific fine-tuning in profile defect identification. Overall, these findings confirm the effectiveness and necessity of all core components within our proposed DGDPO.
Parameter Sensitivity Analysis (RQ4).
Split Ratio $\alpha$ Prec Recall Acc F1
0.5 0.2042 0.7522 0.7540 0.3212
0.6 0.2058 0.7556 0.7570 0.3235
0.7 0.1977 0.7253 0.7472 0.3107
0.8 0.1902 0.7110 0.7411 0.3001
Table 4: The impact of split ratio $\alpha$ on Books dataset.
Specialized LLM Prec Recall Acc F1
Qwen3-0.6B (Ours) 0.2058 0.7556 0.7570 0.3235
Qwen3-1.7B 0.2007 0.7541 0.7522 0.3170
Table 5: The impact of model scale of the specialized LLM in the diagnostic module on Books dataset.
We analyze the impact of split ratio $\alpha$ on the performance of DGDPO. As shown in Table 4, the performance peaks when $\alpha=0.6$ and declines as the ratio deviates from this value. (1) A high $\alpha$ value (e.g., 0.8) leaves limited data for the dynamic optimization, thereby restricting DGDPO’s capacity to detect and correct profile defects, which ultimately degrades the performance. (2) A low $\alpha$ value (e.g., 0.5) results in profile initialization based on less data, increasing the risk of inaccuracy or incompleteness at the outset. Despite the availability of more data for subsequent optimization, the suboptimal starting point constrains final performance. This underscores the necessity of a balanced data allocation strategy to ensure both robust initialization and effective profile optimization.
We also examine the impact of the model scale of the specialized LLM in the diagnostic module. Table 5 compares the performance of our default Qwen3-0.6B model with a larger 1.7B variant. The results show that using the larger 1.7B model does not yield performance gains; instead, the Qwen3-0.6B model performs better across all metrics. A possible reason is that the profile defect identification is a specialized and fine-grained task. Thus, a smaller and more focused model like Qwen3-0.6B can be more effectively and efficiently fine-tuned using our designed training strategy. This finding indicates that for specialized components such as our diagnostic module, model scaling should be guided by task requirements rather than model size alone.
Case Study (RQ5).
Figure 5: The case study on Books dataset.
Fig. 5 provides a concrete illustration of DGDPO’s optimization process on Books dataset. It is triggered by observing a discrepancy case where the simulator built upon $\mathcal{P}_{u}^{(ini)}$ did not choose to interact with the book “Munich, 1938: Appeasement and World War II”, while the real user read it and gave it a high rating of 5. At Step 2, our diagnostic module then correctly identifies this gap and diagnoses the profile as “Incomplete” (in red). Guided by this diagnosis, at Step 3, the treatment module infers the underlying reason (in yellow), and generates a targeted suggestion to add the missing preference for “Politics & Social Sciences” genres (in green) at Step 4. At Step 5, this suggestion is applied to obtain the updated profile $\tilde{\mathcal{P}}_{u}^{(ini)}$ (in blue), making it more comprehensive and capable of explaining the user’s interest in the book. Additional case studies regarding “Inaccurate” and “Inaccurate & Incomplete” are in Appendix E.
Conclusion
In this paper, we propose DGDPO, a novel framework that constructs user profile through a dynamic optimization process to enhance simulation fidelity. DGDPO consists of two core modules: a specialized LLM-based diagnostic module calibrated through our novel training strategy first accurately identifies profile defects, which then guides a generalized LLM-based treatment module to generate targeted refinements. Moreover, we propose to incorporate DGDPO with sequential recommenders to enable realistic multi-round interactions, establishing a bidirectional evolution process between user profile and recommendation strategies. Experiments on three real-world datasets demonstrate the superiority of our DGDPO against the state-of-the-art methods.
Appendix A Data Construction for the Specialized LLM-based Diagnostic Module
A.1 Domain-Adaptive Pre-training Data
We follow the method described in the “Domain-Adaptive Pre-training” section to construct the pre-training data. Each data instance is subsequently converted into a structured text format to effectively pre-train a small-scale general-purpose LLM for injecting recommendation domain-specific semantic and reasoning information. To this end, we design a specific template to integrate the information within each data instance, including the simulated user profile $\mathcal{P}_{u}^{(ini)}$, relevant information from $i_{t}$, and the simulator’s output. The general format of this template is presented in Fig. 6. Then, to specifically illustrate an example of this template, Fig. 7 presents a real training sample derived from the Books dataset.
Figure 6: The structured data template for domain-adaptive pre-training.
Figure 7: An example of a pre-training data sample from the Books dataset.
A.2 Defect-Specific Fine-tuning Data
We follow the data synthesis method described in the “Defect-Specific Fine-tuning” section to synthesize samples for model fine-tuning. All synthetic data are structured into a unified four-part instruction fine-tuning format, as illustrated in Fig. 8. This format clearly defines the System, Instruction, Input, and Output.
Fig. 9 presents specific training samples for three defect types using examples from the Books dataset. To clearly reveal the synthesis method for each defect type, the figure also provides a comparison between the synthesized defective profiles and their original profiles. First, Fig. 9 (a) shows the task-universal System and Instruction prompt. They define the profile defect identification tasks that the LLM needs to perform. This content remains consistent across all synthesized samples. Subsequently, Fig. 9 (b) illustrates the construction of the “Inaccurate” sample. Since the recommended item “The Go-Between” belongs to the “Literature & Fiction” category, we specifically modify the related “Fiction Enthusiast” preference in the original profile, replacing it with an emotionally opposite description “Averse to fiction that lacks depth and originality”. Similarly, Fig. 9 (c) constructs the “Incomplete” sample by directly removing the key preference “Fiction Enthusiast”, rendering the profile information insufficient to explain why the user would favor this item. Finally, Fig. 9 (d) combines both modification approaches to construct the “Inaccurate & Incomplete” sample: it not only removes the “Fiction Enthusiast” preference but simultaneously modifies another preference “Science Fiction Buff” to the emotionally contradictory “Detest predictable fiction stories Buff”.
Figure 8: The four-part instruction-tuning data template for the defect-specific fine-tuning.
Figure 9: Examples of fine-tuning data samples from the Books dataset.
Appendix B Prompts for the Generalized LLM-based Treatment Module
Figure 10: The prompts used in the generalized LLM-based treatment module.
This section provides a detailed presentation of the complete prompt engineering employed by the generalized LLM-based treatment module. This module optimizes user profile through a carefully designed three-step pipeline, which sequentially comprises three core components: first, inferring the defect reasons of the current user profile $\mathcal{P}_{u}^{(ini)}$ based on the diagnostic module’s results; second, generating optimization suggestions based on these reasons; and finally, applying the suggestions to update and optimize the $\mathcal{P}_{u}^{(ini)}$. Fig. 10 comprehensively demonstrates the complete prompts corresponding to these three steps across its three subfigures. It can be observed that each prompt explicitly specifies the LLM’s role, specific task instructions and constraints, as well as structured input-output formats, thereby ensuring the high quality of the entire optimization process.
Appendix C Implementation Details
All experiments are conducted on a server equipped with one NVIDIA V100-SXM2-32GB GPU, two Intel Xeon Platinum 8268 CPUs, and 100GB of memory. In our DGDPO framework, the diagnostic module is based on the Qwen3-0.6B model and trained using the efficient LLaMA-Factory framework [^3] https://github.com/hiyouga/LLaMA-Factory. Specifically, the learning rate is set to 2e-5, batch size to 8, and AdamW is employed as the optimizer. During the pre-training phase, the maximum sequence length is set to 256; in the fine-tuning phase, the maximum sequence length is set to 512. The treatment module utilizes the GPT-4o-mini model through API calls, with the temperature parameter set to 0. For the implementation of the baseline model, we uniformly adopt the code provided by the original authors. Among these, the LLM-based baseline models also invoke GPT-4o-mini (temperature = 0). For detailed hyper-parameter settings of the baseline models, please refer to Table 6.
Appendix D Profile Defect Identification Experiments
We first employ the data synthesis pipeline from the “Defect-Specific Fine-tuning” phase to construct a dataset containing high-quality, labeled defect samples. This dataset is then randomly partitioned into training and testing sets at an 8:2 ratio. The training set is utilized to fine-tune our specialized LLM-based diagnostic module, while the testing set serves to evaluate the performance of both the prompt-based general-purpose LLM and the specialized LLM-based diagnostic module on the defect identification task.
To ensure a fair comparison, we adopt the identical instruction format used for fine-tuning the specialized LLM-based diagnostic module when evaluating the general-purpose LLM method. Specifically, we employ the System and Instruction prompts presented in Fig. 9 (a), combined with the Input component from test samples (following the structure illustrated in Fig. 9(b)-9(d)) to construct complete evaluation prompts. During the evaluation process, the model receives only these three components of information and is required to predict the corresponding defect type. We adopt accuracy as the primary evaluation metric for this experiment, which is calculated as $Accuracy=\frac{N_{\text{correct}}}{N_{\text{total}}}$, where $N_{\text{correct}}$ represents the number of samples in the test set for which the LLM’s predicted defect type matches the ground truth label, and $N_{\text{total}}$ denotes the total number of samples in the test set.
Hyper-parameter ML-1M Books Movies & TV Search Space
KuaiSim embedding_size $32$ $64$ $32$ {32,64,12832,64,128}
learning_rate $10^{-3}$ $10^{-3}$ $10^{-3}$ $10^{-4},10^{-3},10^{-2}$}
batch_size $128$ $128$ $128$ {64,128,25664,128,256}
FPMC embedding_size $64$ $128$ $128$ {32,64,12832,64,128}
learning_rate $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-4},10^{-3},10^{-2}$}
batch_size $128$ $128$ $64$ {64,128,25664,128,256}
NARM embedding_size $64$ $128$ $128$ {32,64,12832,64,128}
learning_rate $10^{-3}$ $10^{-3}$ $10^{-2}$ $10^{-4},10^{-3},10^{-2}$}
batch_size $256$ $128$ $128$ {64,128,25664,128,256}
hidden_size 50 100 100 {50,100,150,20050,100,150,200}
layers 2 3 2 {1,2,31,2,3}
SASRec embedding_size $64$ $32$ $64$ {32,64,12832,64,128}
learning_rate $10^{-2}$ $10^{-3}$ $10^{-3}$ $10^{-4},10^{-3},10^{-2}$}
batch_size $64$ $128$ $128$ {64,128,25664,128,256}
self_attention_head_num 22 11 22 {1,2,4,81,2,4,8}
self_attention_block_num 11 11 11 {1,2,3,41,2,3,4}
LRURec embedding_size $64$ $64$ $64$ {32,64,12832,64,128}
learning_rate $10^{-2}$ $10^{-3}$ $10^{-2}$ $10^{-4},10^{-3},10^{-2}$}
batch_size $128$ $256$ $128$ {64,128,25664,128,256}
Table 6: The search space of hyper-parameters and the optimal settings found by Optuna (optuna.org) for KuaiSim and SRs baselines on the three real-world datasets.
Figure 11: Case studies of the DGDPO optimization process.
Appendix E Additional Case Studies
Fig. 11 provides two additional cases that further demonstrate the DGDPO framework’s capability to handle two other defect types.
Fig. 11 (a) presents an “Inaccurate” case triggered by a direct contradiction: the simulated user profile explicitly states “Disinterested in Folk Tales & Myths”, whereas the real user assigned a high rating (5/5 stars) to “Sleeping Beauty”, a book belonging to this genre. At Step 2, the diagnostic module accurately identifies this contradiction and diagnoses the profile as “Inaccurate” (in red). Guided by this diagnosis, at Step 3, the treatment module determines the cause to be a discrepancy between the profile’s stated preference and the user’s actual behavior (in yellow). Subsequently, at Step 4, the treatment module generates an optimization suggestion to “explicitly state an interest in Folk Tales & Myths” (in green). Finally, at Step 5, this suggestion is implemented to update the user profile, correcting the description to “Have a high interest in Folk Tales & Myths” (in blue), thereby resolving the original contradiction.
Fig. 11 (b) shows an “Inaccurate & Incomplete” case. For the mystery novel “How Far is Heaven?: A Louise Golden Mystery”, which the real user rates 5/5 stars, the defect manifests in two distinct ways: (1)Inaccurate”: the user profile indicates only “moderate interest” in the mystery genre, a mild preference level that fails to align with the user’s demonstrated high enthusiasm (5/5 stars). (2)Incomplete”: the user profile fails to mention a key attribute of this book: “character-driven mysteries”. Given these two issues, at Step 2, the diagnostic module accurately categorizes the defect as “Inaccurate & Incomplete” (in red). At Step 3, the treatment module’s analysis identifies both the aforementioned interest level mismatch (in gray) and the missing preference attribute (in yellow). Correspondingly, at Step 4, the treatment module provides a targeted optimization suggestion to “add a specific preference for ‘character-driven mysteries’” (in green). Finally, at Step 5, the updated profile (in blue) incorporates this suggestion, resulting in a more comprehensive description that better reflects the user’s true preferences.

### v4

@@
 Extensive experiments conducted on three real-world datasets demonstrate the effectiveness of our proposed framework.
-00footnotetext: denotes the corresponding authors.
 Introduction
-User simulators (Balog and Zhai 2023), which aim to mimic user behaviors and decision-making patterns, are crucial for advancing recommender systems (RSs) (Luo et al.
-2022).
+User simulators (Balog:2024:FnTIR), which aim to mimic user behaviors and decision-making patterns, are crucial for advancing recommender systems (RSs) (luo2022mindsim).
 They offer a controllable and cost-effective paradigm for developing and evaluating RSs, reducing the reliance on real-user experiments that are often expensive and ethically sensitive.
-While early user simulators ranged from rule-based (Rohde et al.
-2018) to deep learning-based approaches (Shi et al.
-2019b; Zhao et al.
-2021), these traditional methods fail to explicitly represent users’ complex semantic preferences and generate diverse user interaction behaviors.
-To address these issues, Large Language Models (LLMs) have recently emerged as a new paradigm (Sun et al.
-2024b).
-With their advanced cognitive and generative capabilities, LLMs are typically integrated into an agent framework (Park et al.
-2023) that incorporates components such as Profile, Memory, and Action to emulate user interactions with RSs.
-For example, Agent4Rec (Zhang et al.
-2024a) and RecAgent (Wang et al.
-2025) prompt LLMs to generate user profile, enabling them to simulate diverse behaviors that mirror the complexity of real users.
+While early user simulators ranged from rule-based (rohde2018recogym) to deep learning-based approaches (shi2019virtual; zhao2021usersim), these traditional methods fail to explicitly represent users’ complex semantic preferences and generate diverse user interaction behaviors.
+To address these issues, Large Language Models (LLMs) have recently emerged as a new paradigm.
+With their advanced cognitive and generative capabilities, LLMs are typically integrated into an agent framework (park2023generative) that incorporates components such as Profile, Memory, and Action to emulate user interactions with RSs.
+For example, Agent4Rec (zhang2024generative) and RecAgent (wang2025user) prompt LLMs to generate user profile, enabling them to simulate diverse behaviors that mirror the complexity of real users.
 However, existing LLM-based user simulators face two major limitations: (1) the inaccuracy and incompleteness of profile construction due to the static and single-step prompt-based inference.
@@
 Thus, any inaccuracy or incompleteness in this initial profile persists uncorrected, causing the simulated behavior to progressively diverge from real user actions throughout the simulation; (2) the unrealistic and single-round recommendation-feedback interaction pattern.
-Most existing simulators are typically integrated with static recommenders, e.g., MF (Koren, Bell, and Volinsky 2009), failing to simulate authentic, multi-round interactions where user profile and RS strategies should adapt and evolve based on the latest feedback.
+Most existing simulators are typically integrated with static recommenders, e.g., MF (koren2009matrix), failing to simulate authentic, multi-round interactions where user profile and RS strategies should adapt and evolve based on the latest feedback.
 ML-1MBooksMovies & TV02020404060608080100100Accuracy (%) Prompt-based general-purpose LLM (GPT-4o-mini) Specialized LLM-based diagnostic module (Qwen3-0.6B)
-Figure 1: Accuracy comparison for profile defect identification.
-Our specialized LLM-based diagnostic module significantly outperforms the prompt-based general-purpose LLM.
-Details are included in Appendix D.
+Figure 1: Accuracy comparison for profile defect identification. Our specialized LLM-based diagnostic module significantly outperforms the prompt-based general-purpose LLM.
 Therefore, we propose a novel Diagnostic-Guided Dynamic Profile Optimization (DGDPO) framework for a more realistic simulation.
@@
 (CH1): How to build a reliable mechanism for dynamic profile construction?
-While existing methods (Shu et al.
-2024) like self-reflection (Sun et al.
-2024a) attempt iterative refinement, they often rely on general-purpose LLMs, thus may not be reliable due to potential hallucinations.
+While existing methods (shu2024rah) like self-reflection (sun2024large) attempt iterative refinement, they often rely on general-purpose LLMs, thus may not be reliable due to potential hallucinations.
 As shown in Fig.
@@
 (CH2): How to achieve realistic multi-round interactions where user profiles and recommender strategies mutually influence and evolve over time?
-Inspired by the multi-round interaction nature of real-world recommendation scenarios, we propose to incorporate the user simulator with sequential recommenders (SRs) (Sun et al.
-2025).
+Inspired by the multi-round interaction nature of real-world recommendation scenarios, we propose to incorporate the user simulator with sequential recommenders (SRs) (sun2025llm4rsr).
 This establishes a genuine bidirectional evolution: the user profile is updated based on interactions, while the SRs adaptively adjust their recommendation strategies in response to the evolving user behavior.
@@
 Traditional User Simulators for RSs.
-Early user simulators for RSs progressed from rule-based approaches (Rohde et al.
-2018; Ie et al.
-2019; Shi et al.
-2019a) to advanced data-driven techniques (Huang et al.
-2020), including methods based on reinforcement learning (RL) (Shi et al.
-2019b; Zou et al.
-2020), generative adversarial networks (GANs) (Chen et al.
-2019; Shi et al.
-2019b; Zhao et al.
-2021), and transformers (Zhao et al.
-2023; Afzali et al.
-2023).
+Early user simulators for RSs progressed from rule-based approaches (rohde2018recogym; ie2019recsim; shi2019pyrecgym) to advanced data-driven techniques (huang2020keeping), including methods based on reinforcement learning (RL) (shi2019virtual; zou2020pseudo), generative adversarial networks (GANs) (chen2019generative; shi2019virtual; zhao2021usersim), and transformers (zhao2023kuaisim; afzali2023usersimcrs).
 However, they lack external context and knowledge, as well as the sophisticated cognitive reasoning mechanism that real users leverage during decision-making processes.
 LLM-based User Simulators for RSs.
-The advent of LLMs has introduced novel paradigms for user simulation in RSs (Cai et al.
-2025; Ma et al.
-2025).
-LLM-based simulators have been widely adopted across various applications, including evaluating multi-round conversational systems (Friedman et al.
-2023; Zhu, Huang, and Sang 2024) and generating high-quality interaction data for RL-based algorithms (Corecco et al.
-2024; Ebrat, Paradalis, and Rueda 2024; Zhang et al.
-2025).
+The advent of LLMs has introduced novel paradigms for user simulation in RSs (cai2025agentic; ma2025pub; wei2025mirroring).
+LLM-based simulators have been widely adopted across various applications, including evaluating conversational systems (friedman2023leveraging; zhu2024reliable; zhu2025llm) and generating high-quality interaction data for RL-based algorithms (corecco2024llm; ebrat2024lusifer; zhang2025llm).
 Current works on user profile construction can be categorized into two types.
-Some works (Zhang et al.
-2024a; Wang et al.
-2025) create user profile once at initialization, resulting in static representations that cannot capture evolving user preferences.
-Others employ self-reflection (Zhang et al.
-2024b; Shu et al.
-2024) for iterative refinement.
+Some works (zhang2024generative; wang2025user) create user profile once at initialization, resulting in static representations that cannot capture evolving user preferences.
+Others employ self-reflection (zhang2024agentcf; shu2024rah) for iterative refinement.
 However, they suffer from two limitations: (1) they adopt general-purpose LLMs which may generate unreliable profile due to potential hallucinations; (2) they rely on unrealistic single-round interactions, lacking dynamic adaptation to user feedback and failing to simulate the bidirectional evolution between user profile and recommendation strategies in realistic interactions.
 Sequential Recommenders.
-SRs (Du et al.
-2024) have evolved from early approaches combining Markov Chains with matrix factorization (Rendle, Freudenthaler, and Schmidt-Thieme 2010), to deep learning models.
-These include CNNs for local pattern extraction (Tang and Wang 2018), RNNs to capture temporal dependencies (Hidasi 2015; Huang et al.
-2025), attention mechanisms to focus on relevant historical items (Li et al.
-2017; Kang and McAuley 2018; Liu et al.
-2018; Sun et al.
-2019), and GNNs to model high-order relationships within sequential data (Wu et al.
-2022; Yu et al.
-2024; Li et al.
-2024; Liu, Xia, and Huang 2024).
+SRs have evolved from early approaches combining Markov Chains with matrix factorization (rendle2010factorizing), to deep learning models.
+These include CNNs for local pattern extraction (tang2018personalized), RNNs to capture temporal dependencies (hidasi2015session; zhu2017next; zhu2019query), attention mechanisms to focus on relevant historical items (li2017neural; kang2018self; liu2018stamp; sun2019bert4rec), and GNNs to model high-order relationships within sequential data (wu2022gcrec; yu2024learning; li2024global; liu2024selfgnn).
 Despite their success, they have primarily been evaluated in a static mode, which may fail to fully reflect the dynamic complexity of real-world user interactions.
@@
 For effective and efficient user profile construction, we first create an initial user profile $\mathcal{P}_{u}^{(ini)}$ for each simulated user $u$.
-It is derived from $\mathcal{D}_{u}^{(ini)}$ via a single-step prompt-based inference method (Zhang et al.
-2024a), serving as the foundation user profile.
+It is derived from $\mathcal{D}_{u}^{(ini)}$ via a single-step prompt-based inference method (zhang2024generative), serving as the foundation user profile.
 Then, we leverage $\mathcal{D}_{u}^{(opt)}$ to dynamically optimize $\mathcal{P}_{u}^{(ini)}$ via our proposed DGDPO, ultimately obtaining an optimized user profile $\mathcal{P}_{u}^{(opt)}$.
@@
 To ensure the data quality, we only consider the case where the user simulator’s decision matches with the real user $u$’s behavior (i.e., choosing to interact with $i_{t}$) and user $u$ gave item $i_{t}$ a high rating ($\geq$ 3 out of 5).
-Then, the user simulator’s output, together with $\mathcal{P}_{u}^{(ini)}$ and relevant information (e.g., title, genres, and rating) of $i_{t}$, is used to construct a comprehensive corpus of structured texts, where the details are shown in Appendix A.
+Then, the user simulator’s output, together with $\mathcal{P}_{u}^{(ini)}$ and relevant information (e.g., title, genres, and rating) of $i_{t}$, is used to construct a comprehensive corpus of structured texts.
 Given the constructed corpus, we perform domain-adaptive pre-training on a small-scale general-purpose LLM using a language modeling objective, where the goal is to predict the next token in a sequence.
@@
 (3) Inaccurate & Incomplete: we first apply the “incomplete” modification, followed by the “inaccurate” modification.
-Using these synthetic profiles, we then format the data for instruction fine-tuning, where each sample has four parts: System, Instruction, Input, and Output, as shown in Appendix A.
+Using these synthetic profiles, we then format the data for instruction fine-tuning, where each sample has four parts: System, Instruction, Input, and Output.
 To fine-tune the model, we employ a next token prediction loss objective with a modification: the loss is only calculated on the tokens in the Output field.
@@
 $\mathcal{C}$ $\leftarrow$ [ ];
-  // Initialize discrepancy cases list
+// Initialize
 […]
-  // Single-step prompt-based inference
+// Inference
 1 for item $i$ in $\mathcal{D}_{u}^{(opt)}$ do
 […]
-    // Generate simulated behavior
+ // Get simulated behavior
 […]
@@
 4    if $|\mathcal{C}|=N_{1}$ then
-         // The Specialized LLM-based Diagnostic Module
+       // The Diagnostic Module
 5       $\{l\}\leftarrow f_{\text{diag}}(\mathcal{P}_{u}^{(ini)},\mathcal{C})$;
-         // The Generalized LLM-based Treatment Module
-       $\{r\}\leftarrow\text{Infer}(\mathcal{P}_{u}^{(ini)},\mathcal{C},\{l\})$;
-          // Infer defect reasons
-         // Generate optimization suggestions
+       // The Treatment Module
+6       $\{r\}\leftarrow\text{Infer}(\mathcal{P}_{u}^{(ini)},\mathcal{C},\{l\})$
 7       $\mathcal{O}\leftarrow\text{GenOpt}(\mathcal{P}_{u}^{(ini)},\mathcal{C},\{l\},\{r\})$;
-       $\tilde{\mathcal{P}}_{u}^{(ini)}\leftarrow\text{Refine}(\mathcal{P}_{u}^{(ini)},\mathcal{O})$;
-          // Refine the profile
-7       $\mathcal{C}$ $\leftarrow$ [ ];
+8       $\tilde{\mathcal{P}}_{u}^{(ini)}\leftarrow\text{Refine}(\mathcal{P}_{u}^{(ini)},\mathcal{O})$ $\mathcal{C}$ $\leftarrow$ [ ];
 9       $\mathcal{P}_{u}^{(ini)}$ $\leftarrow$ $\tilde{\mathcal{P}}_{u}^{(ini)}$;
-9
 10
+11
 […]
@@
 While the specialized diagnostic module excels at identifying what defects exist within a profile, understanding why these defects occur and determining how to address them effectively requires sophisticated common-sense reasoning and high-fidelity generative capabilities.
-Consequently, we employ a powerful general-purpose LLM, leveraging its superior analytical reasoning and text generation capabilities to carry out the following treatment process [^1] Detailed prompts are attached in Appendix B..
+Consequently, we employ a powerful general-purpose LLM, leveraging its superior analytical reasoning and text generation capabilities to carry out the following treatment process.
 Specifically, we first prompt the LLM by providing it with the current user profile $\mathcal{P}_{u}^{(ini)}$, the target item $i$, and defect label $l$, to generate a comprehensive textual explanation (i.e., inferred reason $r$) towards the diagnosis made by the diagnosis module.
@@
 Bidirectional Evolution between User Profiles and Recommendation Strategies (CH2)
-Existing LLM-based user simulators are primarily integrated with static recommenders in page-by-page scenarios (Zhang et al.
-2024a), which essentially represent a one-time recommendation result split into multiple pages.
+Existing LLM-based user simulators are primarily integrated with static recommenders in page-by-page scenarios (zhang2024generative), which essentially represent a one-time recommendation result split into multiple pages.
 This fails to capture realistic scenarios where users continuously interact with the RSs, leading to the dynamic evolution of both user profiles and recommendation strategies through user feedback and interactions.
@@
 Experiments and Results
-We conduct extensive experiments to answer five research questions.
+We conduct extensive experiments to answer five research questions [^1] Our code and datasets are available at https://github.com/hyllll/DGDPO.
 (RQ1) Can DGDPO improve the fidelity of user simulators?
@@
 Datasets.
-We use three real-world datasets from various domains.
-Specifically, ML-1M (Harper and Konstan 2015) is widely used in movie recommendations.
-Amazon Books (Books) and Amazon Movies & TV (Movies & TV) are product review datasets collected from Amazon for Books and Movies & TV categories (Ni, Li, and McAuley 2019), respectively.
-We filter out users and items with less than 5 interactions in ML-1M, and those with less than 20 interactions in Amazon datasets.
-To support multi-round interactions, we first sort each user’s interaction chronologically.
-We set the maximum sequence length to 200 for ML-1M and 50 for Amazon datasets.
-For each user $u$, we select the most recent 10 items as the test set, the 11th most recent item as the validation set, and the remaining as the training set (i.e., $\mathcal{D}_{u}$).
-Particularly, we further split $\mathcal{D}_{u}$ into $\mathcal{D}_{u}^{(ini)}$ and $\mathcal{D}_{u}^{(opt)}$ with the ratio of $\alpha:(1-\alpha)$.
-The statistics of these datasets after preprocessing are summarized in Table 1.
+We use three datasets: ML-1M (harper2015movielens), Amazon Books, and Amazon Movies & TV (ni2019justifying).
+We filter users and items with less than 5 interactions (ML-1M) or less than 20 (Amazon datasets).
+Interactions are sorted chronologically, with max sequence lengths set to 200 for ML-1M and 50 for Amazon datasets.
+For each user, the last 10 items are for testing, the 11th most recent item is for validation, and the rest for training (i.e., $\mathcal{D}_{u}$).
+$\mathcal{D}_{u}$ is further split into $\mathcal{D}_{u}^{(ini)}$ and $\mathcal{D}_{u}^{(opt)}$ by an $\alpha:(1-\alpha)$ ratio.
 Baselines.
 We compare DGDPO with five baselines for user profile construction in two categories.
-In traditional methods, KuaiSim (Zhao et al.
-2023) is a transformer-based model for user profile construction.
-In LLM-based methods, Agent4Rec (Zhang et al.
-2024a) utilizes a single-step prompt-based inference.
-LLM-US (Zhang et al.
-2025) generates explanations of user ratings as profile.
-RAH (Shu et al.
-2024) and AgentCF (Zhang et al.
-2024b) employ self-reflection and collaborative reflection, respectively.
-To enable multi-round interactions, we integrate the user simulator with four representative SRs.
-The conventional method FPMC (Rendle, Freudenthaler, and Schmidt-Thieme 2010) combines matrix factorization with first-order Markov chains.
-The deep learning-based methods include NARM (Li et al.
-2017), an RNN-based model with an attention mechanism; SASRec (Kang and McAuley 2018), using unidirectional self-attention; and LRURec (Yue et al.
-2024), leveraging a linear recurrent unit-based model.
+In traditional methods, KuaiSim (zhao2023kuaisim) is a transformer-based model for user profile construction.
+In LLM-based methods, Agent4Rec (zhang2024generative) utilizes a single-step prompt-based inference.
+LLM-US (zhang2025llm) generates explanations of user ratings as profile.
+RAH (shu2024rah) and AgentCF (zhang2024agentcf) employ self-reflection and collaborative reflection, respectively.
+For multi-round interactions, we integrate the user simulator with four SRs: FPMC (rendle2010factorizing), NARM (li2017neural), SASRec (kang2018self) and LRURec (yue2024linear).
 Implementation Details.
-Following state-of-the-arts (Shu et al.
-2024; Zhang et al.
-2024a), we randomly sample 1000 users from each dataset to serve as simulated users.
-Our diagnostic module is based on lightweight Qwen3-0.6B (Yang et al.
-2025), while the treatment module employs gpt-4o-mini via its API with a temperature of 0.
+Following state-of-the-arts (shu2024rah; zhang2024generative), we randomly sample 1000 users from each dataset to serve as simulated users.
+Our diagnostic module is based on lightweight Qwen3-0.6B (yang2025qwen3), while the treatment module employs gpt-4o-mini via its API with a temperature of 0.
 All LLM-based baselines also utilize GPT-4o-mini as their backbone.
 For DGDPO, we set $N_{1}=4$, $\alpha=0.6$, and $\mathcal{X}=20$ for all datasets.
-We build upon Agent4Rec (Zhang et al.
-2024a) as the foundation user simulator and conduct ten-round interactions for each user.
-For KuaiSim and SRs methods, we tune them by Optuna (optuna.org) with 50 trials and train for up to 100 epochs with an early stopping mechanism (Sun et al.
-2022).
-All experiments are conducted on an NVIDIA V100 GPU.
-More detailed implementations are in Appendix C.
+We build upon Agent4Rec (zhang2024generative) as the foundation user simulator and conduct ten-round interactions for each user.
+For KuaiSim and SRs methods, we tune them by Optuna (optuna.org) with 50 trials and train for up to 100 epochs with an early stopping mechanism (sun2022daisyrec).
+All experiments are conducted on a V100 GPU.
 Evaluation Metrics.
-Following (Luo et al.
-2022; Zhang et al.
-2024a), we adopt Precision (Prec), Recall, Accuracy (Acc), and F1 Score (F1) as the evaluation metrics to measure the consistency between the simulated and real users.
+Following (luo2022mindsim; zhang2024generative), we adopt Precision (Prec), Recall, Accuracy (Acc), and F1 Score (F1) as the evaluation metrics to measure the consistency between the simulated and real users.
 Results and Analysis
@@
 In contrast, with advanced deep learning-based SRs, we observe steady performance improvements.
-This occurs because advanced SRs rank ground truth items higher, and LLM-based user simulators, sensitive to item positions (Hou et al.
-2024), are more likely to interact with them when they are prominently placed, resulting in more accurate simulation.
+This occurs because advanced SRs rank ground truth items higher, and LLM-based user simulators, sensitive to item positions (hou2024large), are more likely to interact with them when they are prominently placed, resulting in more accurate simulation.
 (3)Update w/ GT’ performs the best among all methods.
@@
 Fig.
-5 provides a concrete illustration of DGDPO’s optimization process on Books dataset.
-It is triggered by observing a discrepancy case where the simulator built upon $\mathcal{P}_{u}^{(ini)}$ did not choose to interact with the book “Munich, 1938: Appeasement and World War II”, while the real user read it and gave it a high rating of 5.
-At Step 2, our diagnostic module then correctly identifies this gap and diagnoses the profile as “Incomplete” (in red).
-Guided by this diagnosis, at Step 3, the treatment module infers the underlying reason (in yellow), and generates a targeted suggestion to add the missing preference for “Politics & Social Sciences” genres (in green) at Step 4.
-At Step 5, this suggestion is applied to obtain the updated profile $\tilde{\mathcal{P}}_{u}^{(ini)}$ (in blue), making it more comprehensive and capable of explaining the user’s interest in the book.
-Additional case studies regarding “Inaccurate” and “Inaccurate & Incomplete” are in Appendix E.
+5 illustrates the DGDPO optimization process with a case study from the Books dataset.
+A discrepancy (a highly-rated book being ignored by the simulator) triggers the diagnostic module to identify the profile as “Incomplete”.
+The treatment module then infers the missing preference for “Politics & Social Sciences” and updates the profile accordingly.
 Conclusion
@@
 Experiments on three real-world datasets demonstrate the superiority of our DGDPO against the state-of-the-art methods.
-Appendix A Data Construction for the Specialized LLM-based Diagnostic Module
-A.1 Domain-Adaptive Pre-training Data
-We follow the method described in the “Domain-Adaptive Pre-training” section to construct the pre-training data.
-Each data instance is subsequently converted into a structured text format to effectively pre-train a small-scale general-purpose LLM for injecting recommendation domain-specific semantic and reasoning information.
-To this end, we design a specific template to integrate the information within each data instance, including the simulated user profile $\mathcal{P}_{u}^{(ini)}$, relevant information from $i_{t}$, and the simulator’s output.
-The general format of this template is presented in Fig.
-6.
-Then, to specifically illustrate an example of this template, Fig.
-7 presents a real training sample derived from the Books dataset.
-Figure 6: The structured data template for domain-adaptive pre-training.
-Figure 7: An example of a pre-training data sample from the Books dataset.
-A.2 Defect-Specific Fine-tuning Data
-We follow the data synthesis method described in the “Defect-Specific Fine-tuning” section to synthesize samples for model fine-tuning.
-All synthetic data are structured into a unified four-part instruction fine-tuning format, as illustrated in Fig.
-8.
-This format clearly defines the System, Instruction, Input, and Output.
-Fig.
-9 presents specific training samples for three defect types using examples from the Books dataset.
-To clearly reveal the synthesis method for each defect type, the figure also provides a comparison between the synthesized defective profiles and their original profiles.
-First, Fig.
-9 (a) shows the task-universal System and Instruction prompt.
-They define the profile defect identification tasks that the LLM needs to perform.
-This content remains consistent across all synthesized samples.
-Subsequently, Fig.
-9 (b) illustrates the construction of the “Inaccurate” sample.
-Since the recommended item “The Go-Between” belongs to the “Literature & Fiction” category, we specifically modify the related “Fiction Enthusiast” preference in the original profile, replacing it with an emotionally opposite description “Averse to fiction that lacks depth and originality”.
-Similarly, Fig.
-9 (c) constructs the “Incomplete” sample by directly removing the key preference “Fiction Enthusiast”, rendering the profile information insufficient to explain why the user would favor this item.
-Finally, Fig.
-9 (d) combines both modification approaches to construct the “Inaccurate & Incomplete” sample: it not only removes the “Fiction Enthusiast” preference but simultaneously modifies another preference “Science Fiction Buff” to the emotionally contradictory “Detest predictable fiction stories Buff”.
-Figure 8: The four-part instruction-tuning data template for the defect-specific fine-tuning.
-Figure 9: Examples of fine-tuning data samples from the Books dataset.
-Appendix B Prompts for the Generalized LLM-based Treatment Module
-Figure 10: The prompts used in the generalized LLM-based treatment module.
-This section provides a detailed presentation of the complete prompt engineering employed by the generalized LLM-based treatment module.
-This module optimizes user profile through a carefully designed three-step pipeline, which sequentially comprises three core components: first, inferring the defect reasons of the current user profile $\mathcal{P}_{u}^{(ini)}$ based on the diagnostic module’s results; second, generating optimization suggestions based on these reasons; and finally, applying the suggestions to update and optimize the $\mathcal{P}_{u}^{(ini)}$.
-Fig.
-10 comprehensively demonstrates the complete prompts corresponding to these three steps across its three subfigures.
-It can be observed that each prompt explicitly specifies the LLM’s role, specific task instructions and constraints, as well as structured input-output formats, thereby ensuring the high quality of the entire optimization process.
-Appendix C Implementation Details
-All experiments are conducted on a server equipped with one NVIDIA V100-SXM2-32GB GPU, two Intel Xeon Platinum 8268 CPUs, and 100GB of memory.
-In our DGDPO framework, the diagnostic module is based on the Qwen3-0.6B model and trained using the efficient LLaMA-Factory framework [^3] https://github.com/hiyouga/LLaMA-Factory.
-Specifically, the learning rate is set to 2e-5, batch size to 8, and AdamW is employed as the optimizer.
-During the pre-training phase, the maximum sequence length is set to 256; in the fine-tuning phase, the maximum sequence length is set to 512.
-The treatment module utilizes the GPT-4o-mini model through API calls, with the temperature parameter set to 0.
-For the implementation of the baseline model, we uniformly adopt the code provided by the original authors.
-Among these, the LLM-based baseline models also invoke GPT-4o-mini (temperature = 0).
-For detailed hyper-parameter settings of the baseline models, please refer to Table 6.
-Appendix D Profile Defect Identification Experiments
-We first employ the data synthesis pipeline from the “Defect-Specific Fine-tuning” phase to construct a dataset containing high-quality, labeled defect samples.
-This dataset is then randomly partitioned into training and testing sets at an 8:2 ratio.
-The training set is utilized to fine-tune our specialized LLM-based diagnostic module, while the testing set serves to evaluate the performance of both the prompt-based general-purpose LLM and the specialized LLM-based diagnostic module on the defect identification task.
-To ensure a fair comparison, we adopt the identical instruction format used for fine-tuning the specialized LLM-based diagnostic module when evaluating the general-purpose LLM method.
-Specifically, we employ the System and Instruction prompts presented in Fig.
-9 (a), combined with the Input component from test samples (following the structure illustrated in Fig.
-9(b)-9(d)) to construct complete evaluation prompts.
-During the evaluation process, the model receives only these three components of information and is required to predict the corresponding defect type.
-We adopt accuracy as the primary evaluation metric for this experiment, which is calculated as $Accuracy=\frac{N_{\text{correct}}}{N_{\text{total}}}$, where $N_{\text{correct}}$ represents the number of samples in the test set for which the LLM’s predicted defect type matches the ground truth label, and $N_{\text{total}}$ denotes the total number of samples in the test set.
-Hyper-parameter ML-1M Books Movies & TV Search Space
-KuaiSim embedding_size $32$ $64$ $32$ {32,64,12832,64,128}
-learning_rate $10^{-3}$ $10^{-3}$ $10^{-3}$ $10^{-4},10^{-3},10^{-2}$}
-batch_size $128$ $128$ $128$ {64,128,25664,128,256}
-FPMC embedding_size $64$ $128$ $128$ {32,64,12832,64,128}
-learning_rate $10^{-2}$ $10^{-2}$ $10^{-2}$ $10^{-4},10^{-3},10^{-2}$}
-batch_size $128$ $128$ $64$ {64,128,25664,128,256}
-NARM embedding_size $64$ $128$ $128$ {32,64,12832,64,128}
-learning_rate $10^{-3}$ $10^{-3}$ $10^{-2}$ $10^{-4},10^{-3},10^{-2}$}
-batch_size $256$ $128$ $128$ {64,128,25664,128,256}
-hidden_size 50 100 100 {50,100,150,20050,100,150,200}
-layers 2 3 2 {1,2,31,2,3}
-SASRec embedding_size $64$ $32$ $64$ {32,64,12832,64,128}
-learning_rate $10^{-2}$ $10^{-3}$ $10^{-3}$ $10^{-4},10^{-3},10^{-2}$}
-batch_size $64$ $128$ $128$ {64,128,25664,128,256}
-self_attention_head_num 22 11 22 {1,2,4,81,2,4,8}
-self_attention_block_num 11 11 11 {1,2,3,41,2,3,4}
-LRURec embedding_size $64$ $64$ $64$ {32,64,12832,64,128}
-learning_rate $10^{-2}$ $10^{-3}$ $10^{-2}$ $10^{-4},10^{-3},10^{-2}$}
-batch_size $128$ $256$ $128$ {64,128,25664,128,256}
-Table 6: The search space of hyper-parameters and the optimal settings found by Optuna (optuna.org) for KuaiSim and SRs baselines on the three real-world datasets.
-Figure 11: Case studies of the DGDPO optimization process.
-Appendix E Additional Case Studies
-Fig. 11 provides two additional cases that further demonstrate the DGDPO framework’s capability to handle two other defect types.
-Fig.
-11 (a) presents an “Inaccurate” case triggered by a direct contradiction: the simulated user profile explicitly states “Disinterested in Folk Tales & Myths”, whereas the real user assigned a high rating (5/5 stars) to “Sleeping Beauty”, a book belonging to this genre.
-At Step 2, the diagnostic module accurately identifies this contradiction and diagnoses the profile as “Inaccurate” (in red).
-Guided by this diagnosis, at Step 3, the treatment module determines the cause to be a discrepancy between the profile’s stated preference and the user’s actual behavior (in yellow).
-Subsequently, at Step 4, the treatment module generates an optimization suggestion to “explicitly state an interest in Folk Tales & Myths” (in green).
-Finally, at Step 5, this suggestion is implemented to update the user profile, correcting the description to “Have a high interest in Folk Tales & Myths” (in blue), thereby resolving the original contradiction.
-Fig.
-11 (b) shows an “Inaccurate & Incomplete” case.
-For the mystery novel “How Far is Heaven?: A Louise Golden Mystery”, which the real user rates 5/5 stars, the defect manifests in two distinct ways: (1)Inaccurate”: the user profile indicates only “moderate interest” in the mystery genre, a mild preference level that fails to align with the user’s demonstrated high enthusiasm (5/5 stars).
-(2)Incomplete”: the user profile fails to mention a key attribute of this book: “character-driven mysteries”.
-Given these two issues, at Step 2, the diagnostic module accurately categorizes the defect as “Inaccurate & Incomplete” (in red).
-At Step 3, the treatment module’s analysis identifies both the aforementioned interest level mismatch (in gray) and the missing preference attribute (in yellow).
-Correspondingly, at Step 4, the treatment module provides a targeted optimization suggestion to “add a specific preference for ‘character-driven mysteries’” (in green).
-Finally, at Step 5, the updated profile (in blue) incorporates this suggestion, resulting in a more comprehensive description that better reflects the user’s true preferences.
+Acknowledgments
+This research is supported by the Australian Research Council (ARC) Discovery Project DP230100676 and partially supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 1, SUTD Kickstarter Initiative (SKI 2021_06_12).

### v5

@@
 Introduction
-User simulators (Balog:2024:FnTIR), which aim to mimic user behaviors and decision-making patterns, are crucial for advancing recommender systems (RSs) (luo2022mindsim).
+User simulators (Balog and Zhai 2024), which aim to mimic user behaviors and decision-making patterns, are crucial for advancing recommender systems (RSs) (Luo et al.
+2022).
 They offer a controllable and cost-effective paradigm for developing and evaluating RSs, reducing the reliance on real-user experiments that are often expensive and ethically sensitive.
-While early user simulators ranged from rule-based (rohde2018recogym) to deep learning-based approaches (shi2019virtual; zhao2021usersim), these traditional methods fail to explicitly represent users’ complex semantic preferences and generate diverse user interaction behaviors.
+While early user simulators ranged from rule-based (Rohde et al.
+2018) to deep learning-based approaches (Shi et al.
+2019b; Zhao et al.
+2021), these traditional methods fail to explicitly represent users’ complex semantic preferences and generate diverse user interaction behaviors.
 To address these issues, Large Language Models (LLMs) have recently emerged as a new paradigm.
-With their advanced cognitive and generative capabilities, LLMs are typically integrated into an agent framework (park2023generative) that incorporates components such as Profile, Memory, and Action to emulate user interactions with RSs.
-For example, Agent4Rec (zhang2024generative) and RecAgent (wang2025user) prompt LLMs to generate user profile, enabling them to simulate diverse behaviors that mirror the complexity of real users.
+With their advanced cognitive and generative capabilities, LLMs are typically integrated into an agent framework (Park et al.
+2023) that incorporates components such as Profile, Memory, and Action to emulate user interactions with RSs.
+For example, Agent4Rec (Zhang et al.
+2024a) and RecAgent (Wang et al.
+2025) prompt LLMs to generate user profile, enabling them to simulate diverse behaviors that mirror the complexity of real users.
 However, existing LLM-based user simulators face two major limitations: (1) the inaccuracy and incompleteness of profile construction due to the static and single-step prompt-based inference.
@@
 Thus, any inaccuracy or incompleteness in this initial profile persists uncorrected, causing the simulated behavior to progressively diverge from real user actions throughout the simulation; (2) the unrealistic and single-round recommendation-feedback interaction pattern.
-Most existing simulators are typically integrated with static recommenders, e.g., MF (koren2009matrix), failing to simulate authentic, multi-round interactions where user profile and RS strategies should adapt and evolve based on the latest feedback.
+Most existing simulators are typically integrated with static recommenders, e.g., MF (Koren et al.
+2009), failing to simulate authentic, multi-round interactions where user profile and RS strategies should adapt and evolve based on the latest feedback.
 ML-1MBooksMovies & TV02020404060608080100100Accuracy (%) Prompt-based general-purpose LLM (GPT-4o-mini) Specialized LLM-based diagnostic module (Qwen3-0.6B)
@@
 (CH1): How to build a reliable mechanism for dynamic profile construction?
-While existing methods (shu2024rah) like self-reflection (sun2024large) attempt iterative refinement, they often rely on general-purpose LLMs, thus may not be reliable due to potential hallucinations.
+While existing methods (Shu et al.
+2024) like self-reflection (Sun et al.
+2024) attempt iterative refinement, they often rely on general-purpose LLMs, thus may not be reliable due to potential hallucinations.
 As shown in Fig.
@@
 (CH2): How to achieve realistic multi-round interactions where user profiles and recommender strategies mutually influence and evolve over time?
-Inspired by the multi-round interaction nature of real-world recommendation scenarios, we propose to incorporate the user simulator with sequential recommenders (SRs) (sun2025llm4rsr).
+Inspired by the multi-round interaction nature of real-world recommendation scenarios, we propose to incorporate the user simulator with sequential recommenders (SRs) (Sun et al.
+2025).
 This establishes a genuine bidirectional evolution: the user profile is updated based on interactions, while the SRs adaptively adjust their recommendation strategies in response to the evolving user behavior.
@@
 Traditional User Simulators for RSs.
-Early user simulators for RSs progressed from rule-based approaches (rohde2018recogym; ie2019recsim; shi2019pyrecgym) to advanced data-driven techniques (huang2020keeping), including methods based on reinforcement learning (RL) (shi2019virtual; zou2020pseudo), generative adversarial networks (GANs) (chen2019generative; shi2019virtual; zhao2021usersim), and transformers (zhao2023kuaisim; afzali2023usersimcrs).
+Early user simulators for RSs progressed from rule-based approaches (Rohde et al.
+2018; Ie et al.
+2019; Shi et al.
+2019a) to advanced data-driven techniques (Huang et al.
+2020), including methods based on reinforcement learning (RL) (Shi et al.
+2019b; Zou et al.
+2020), generative adversarial networks (GANs) (Chen et al.
+2019; Shi et al.
+2019b; Zhao et al.
+2021), and transformers (Zhao et al.
+2023; Afzali et al.
+2023).
 However, they lack external context and knowledge, as well as the sophisticated cognitive reasoning mechanism that real users leverage during decision-making processes.
 LLM-based User Simulators for RSs.
-The advent of LLMs has introduced novel paradigms for user simulation in RSs (cai2025agentic; ma2025pub; wei2025mirroring).
-LLM-based simulators have been widely adopted across various applications, including evaluating conversational systems (friedman2023leveraging; zhu2024reliable; zhu2025llm) and generating high-quality interaction data for RL-based algorithms (corecco2024llm; ebrat2024lusifer; zhang2025llm).
+The advent of LLMs has introduced novel paradigms for user simulation in RSs (Cai et al.
+2025; Ma et al.
+2025; Wei et al.
+2025).
+LLM-based simulators have been widely adopted across various applications, including evaluating conversational systems (Friedman et al.
+2023; Zhu et al.
+2024, 2025) and generating high-quality interaction data for RL-based algorithms (Corecco et al.
+2024; Ebrat et al.
+2024; Zhang et al.
+2025).
 Current works on user profile construction can be categorized into two types.
-Some works (zhang2024generative; wang2025user) create user profile once at initialization, resulting in static representations that cannot capture evolving user preferences.
-Others employ self-reflection (zhang2024agentcf; shu2024rah) for iterative refinement.
+Some works (Zhang et al.
+2024a; Wang et al.
+2025) create user profile once at initialization, resulting in static representations that cannot capture evolving user preferences.
+Others employ self-reflection (Zhang et al.
+2024b; Shu et al.
+2024) for iterative refinement.
 However, they suffer from two limitations: (1) they adopt general-purpose LLMs which may generate unreliable profile due to potential hallucinations; (2) they rely on unrealistic single-round interactions, lacking dynamic adaptation to user feedback and failing to simulate the bidirectional evolution between user profile and recommendation strategies in realistic interactions.
 Sequential Recommenders.
-SRs have evolved from early approaches combining Markov Chains with matrix factorization (rendle2010factorizing), to deep learning models.
-These include CNNs for local pattern extraction (tang2018personalized), RNNs to capture temporal dependencies (hidasi2015session; zhu2017next; zhu2019query), attention mechanisms to focus on relevant historical items (li2017neural; kang2018self; liu2018stamp; sun2019bert4rec), and GNNs to model high-order relationships within sequential data (wu2022gcrec; yu2024learning; li2024global; liu2024selfgnn).
+SRs have evolved from early approaches combining Markov Chains with matrix factorization (Rendle et al.
+2010), to deep learning models.
+These include CNNs for local pattern extraction (Tang and Wang 2018), RNNs to capture temporal dependencies (Hidasi 2015; Zhu et al.
+2017, 2019), attention mechanisms to focus on relevant historical items (Li et al.
+2017; Kang and McAuley 2018; Liu et al.
+2018; Sun et al.
+2019), and GNNs to model high-order relationships within sequential data (Wu et al.
+2022; Yu et al.
+2024; Li et al.
+2024; Liu et al.
+2024).
 Despite their success, they have primarily been evaluated in a static mode, which may fail to fully reflect the dynamic complexity of real-world user interactions.
@@
 For effective and efficient user profile construction, we first create an initial user profile $\mathcal{P}_{u}^{(ini)}$ for each simulated user $u$.
-It is derived from $\mathcal{D}_{u}^{(ini)}$ via a single-step prompt-based inference method (zhang2024generative), serving as the foundation user profile.
+It is derived from $\mathcal{D}_{u}^{(ini)}$ via a single-step prompt-based inference method (Zhang et al.
+2024a), serving as the foundation user profile.
 Then, we leverage $\mathcal{D}_{u}^{(opt)}$ to dynamically optimize $\mathcal{P}_{u}^{(ini)}$ via our proposed DGDPO, ultimately obtaining an optimized user profile $\mathcal{P}_{u}^{(opt)}$.
@@
 Bidirectional Evolution between User Profiles and Recommendation Strategies (CH2)
-Existing LLM-based user simulators are primarily integrated with static recommenders in page-by-page scenarios (zhang2024generative), which essentially represent a one-time recommendation result split into multiple pages.
+Existing LLM-based user simulators are primarily integrated with static recommenders in page-by-page scenarios (Zhang et al.
+2024a), which essentially represent a one-time recommendation result split into multiple pages.
 This fails to capture realistic scenarios where users continuously interact with the RSs, leading to the dynamic evolution of both user profiles and recommendation strategies through user feedback and interactions.
@@
 Datasets.
-We use three datasets: ML-1M (harper2015movielens), Amazon Books, and Amazon Movies & TV (ni2019justifying).
+We use three datasets: ML-1M (Harper and Konstan 2015), Amazon Books, and Amazon Movies & TV (Ni et al.
+2019).
 We filter users and items with less than 5 interactions (ML-1M) or less than 20 (Amazon datasets).
@@
 We compare DGDPO with five baselines for user profile construction in two categories.
-In traditional methods, KuaiSim (zhao2023kuaisim) is a transformer-based model for user profile construction.
-In LLM-based methods, Agent4Rec (zhang2024generative) utilizes a single-step prompt-based inference.
-LLM-US (zhang2025llm) generates explanations of user ratings as profile.
-RAH (shu2024rah) and AgentCF (zhang2024agentcf) employ self-reflection and collaborative reflection, respectively.
-For multi-round interactions, we integrate the user simulator with four SRs: FPMC (rendle2010factorizing), NARM (li2017neural), SASRec (kang2018self) and LRURec (yue2024linear).
+In traditional methods, KuaiSim (Zhao et al.
+2023) is a transformer-based model for user profile construction.
+In LLM-based methods, Agent4Rec (Zhang et al.
+2024a) utilizes a single-step prompt-based inference.
+LLM-US (Zhang et al.
+2025) generates explanations of user ratings as profile.
+RAH (Shu et al.
+2024) and AgentCF (Zhang et al.
+2024b) employ self-reflection and collaborative reflection, respectively.
+For multi-round interactions, we integrate the user simulator with four SRs: FPMC (Rendle et al.
+2010), NARM (Li et al.
+2017), SASRec (Kang and McAuley 2018) and LRURec (Yue et al.
+2024).
 Implementation Details.
-Following state-of-the-arts (shu2024rah; zhang2024generative), we randomly sample 1000 users from each dataset to serve as simulated users.
-Our diagnostic module is based on lightweight Qwen3-0.6B (yang2025qwen3), while the treatment module employs gpt-4o-mini via its API with a temperature of 0.
+Following state-of-the-arts (Shu et al.
+2024; Zhang et al.
+2024a), we randomly sample 1000 users from each dataset to serve as simulated users.
+Our diagnostic module is based on lightweight Qwen3-0.6B (Yang et al.
+2025), while the treatment module employs gpt-4o-mini via its API with a temperature of 0.
 All LLM-based baselines also utilize GPT-4o-mini as their backbone.
 For DGDPO, we set $N_{1}=4$, $\alpha=0.6$, and $\mathcal{X}=20$ for all datasets.
-We build upon Agent4Rec (zhang2024generative) as the foundation user simulator and conduct ten-round interactions for each user.
-For KuaiSim and SRs methods, we tune them by Optuna (optuna.org) with 50 trials and train for up to 100 epochs with an early stopping mechanism (sun2022daisyrec).
+We build upon Agent4Rec (Zhang et al.
+2024a) as the foundation user simulator and conduct ten-round interactions for each user.
+For KuaiSim and SRs methods, we tune them by Optuna (optuna.org) with 50 trials and train for up to 100 epochs with an early stopping mechanism (Sun et al.
+2022).
 All experiments are conducted on a V100 GPU.
 Evaluation Metrics.
-Following (luo2022mindsim; zhang2024generative), we adopt Precision (Prec), Recall, Accuracy (Acc), and F1 Score (F1) as the evaluation metrics to measure the consistency between the simulated and real users.
+Following (Luo et al.
+2022; Zhang et al.
+2024a), we adopt Precision (Prec), Recall, Accuracy (Acc), and F1 Score (F1) as the evaluation metrics to measure the consistency between the simulated and real users.
 Results and Analysis
@@
 In contrast, with advanced deep learning-based SRs, we observe steady performance improvements.
-This occurs because advanced SRs rank ground truth items higher, and LLM-based user simulators, sensitive to item positions (hou2024large), are more likely to interact with them when they are prominently placed, resulting in more accurate simulation.
+This occurs because advanced SRs rank ground truth items higher, and LLM-based user simulators, sensitive to item positions (Hou et al.
+2024), are more likely to interact with them when they are prominently placed, resulting in more accurate simulation.
 (3)Update w/ GT’ performs the best among all methods.